### v1

Abstract
Semantic-based test generators are widely used to produce failure-inducing inputs for Deep Learning (DL) systems. They typically generate challenging test inputs by applying random perturbations to input semantic concepts until a failure is found or a timeout is reached. However, such randomness may hinder them from efficiently achieving their goal. This paper proposes XMutant, a technique that leverages explainable artificial intelligence (XAI) techniques to generate challenging test inputs. XMutant uses the local explanation of the input to inform the fuzz testing process and effectively guide it toward failures of the DL system under test. We evaluated different configurations of XMutant in triggering failures for different DL systems both for model-level (sentiment analysis, digit recognition) and system-level testing (advanced driving assistance). Our studies showed that XMutant enables more effective and efficient test generation by focusing on the most impactful parts of the input. XMutant generates up to $125\%$ more failure-inducing inputs compared to an existing baseline, up to 7$\times$ faster. We also assessed the validity of these inputs, maintaining a validation rate above $89\%$, according to automated and human validators.
journal: Journal
Keywords. Software testing, testing deep learning systems, explainable AI
1 Introduction
Deep Learning (DL) systems play a crucial role in software engineering schutze2008introduction due to their ability to solve complex tasks by learning from a corpus of data. However, DL systems pose additional challenges to testing. Unlike traditional software, where behavior can be anticipated through source code analysis, the actions of DL systems are less predictable 2020-Riccio-EMSE . Therefore, test generation for DL systems is critical for ensuring their reliability and correctness. This process involves creating a diverse set of input data to uncover potential weaknesses or biases in the model, such as incorrect predictions or failures to generalize from data samples that are different from those available during training. Testing is particularly important in applications where DL is used for decision-making in critical scenarios, such as autonomous driving 2020-Stocco-ICSE ; Abdessalem-ICSE18 ; Gambi:2019:ATS:3293882.3330566 ; survey-lei-ma . Exhaustive testing of DL systems is impractical because of the size of the input space. Thus, generating effective test cases for these systems is an important problem, which requires a comprehensive understanding of the DL model’s architecture, its training data, and the application domain.
Prior work proposed solutions to generate test cases automatically for DL systems 2020-Riccio-FSE ; Abdessalem-ASE18-1 ; Abdessalem-ASE18-2 ; Abdessalem-ICSE18 ; 9712397 ; Guannam_ETAL_FSE22 ; MULLINS2018197 ; Gambi:2019:ATS:3293882.3330566 ; drivefuzz ; zhongETAL2021 ; liETAL2020 ; Jha2019MLBasedFI ; 10.1145/3597926.3598072 . Some approaches such as DeepXplore pei2017deepxplore , DLFuzz guo2018dlfuzz and DeepTest deeptest target DL image classification systems and they involve raw input manipulation techniques that modify/corrupt pixel values of an input. These techniques do not generate new functional inputs as they produce minimal, often imperceptible changes to the original inputs, and are therefore suitable to test robustness and security deficiencies of the DL system 2023-Riccio-ICSE ; 2025-Maryam-ICST . In contrast, functional test generation focuses on creating new inputs that deviate significantly from the original training distribution. These inputs target the long-tail problem of DL testing zhang2024systematicreviewlongtailedlearning , testing the DNN’s ability to generalize to novel, unseen scenarios. Instances of functional test generators are the model-based approaches (also called semantic-based approaches) like DeepJanus 2020-Riccio-FSE , DeepHyperion zohdinasab2021deephyperion and DeepMetis 2021-Riccio-ASE or latent space manipulation techniques like SINVAD kang2020sinvad ; sinvad-tosem and CIT4DNN dola2024cit4dnn . They generate new inputs with mutations that are randomly applied to a semantic representation of the inputs 2020-Riccio-FSE , or to a latent vector kang2020sinvad . While random mutations can eventually produce inputs that expose failures, these solutions are extraneous to the internal state of the DL system. Indeed, finding failures with black-box approaches is time-consuming, especially in computationally demanding contexts, such as simulation-based testing of self-driving cars.
This paper investigates the development of more effective and efficient semantic-based test generators for DL systems, leveraging eXplainable Artificial Intelligence (XAI) techniques. While methods similar to XAI, leveraging neuron activation values and gradient information, have been explored for raw input manipulation neural_coverage_new ; Ma:2018:DMT:3238147.3238202 ; Xie-ISSTA-2019 to derive adversarial attacks for robustness testing, their application to semantic-based functional test generation remains unexplored. Therefore, we focus on semantic-based approaches that utilize an input model, as they have been successfully applied to various DL systems (e.g., classification, regression) and input types (e.g., text, images, logical driving scenarios).
In this work, we leverage the post-hoc local explanations of DL systems for individual predictions. Local explanations reveal the contributions of features in the input with respect to the model prediction du2019techniques . Depending on the model’s input, the explanations can take the form of feature contributions (e.g., for textual inputs), or visual heatmaps (e.g., for image inputs) jetley2018learn ; samek2019explainable ; samek2017explainable . Previous work highlights the valuable insights provided by XAI, particularly for understanding and debugging DL systems abs-2201-00009 ; 2022-Stocco-ASE ; arxiv.2204.00480 ; DBLP:journals/corr/abs-2002-00863 . In our work, we leverage their information for guiding test generation by introducing XMutant, a DL testing technique that leverages XAI’s explanations to derive challenging inputs. More specifically, XMutant leverages a novel mutation operator that uses the local explanation on a given input to identify the area of its semantic representation that has higher contributions to the decision-making process of DL systems. Our work shows that fuzzing such attention areas with targeted minor modifications accelerates fault exposure and preserves validity (i.e., the realism of the inputs) and label (i.e., the oracle). XMutant uses the local explanations in two ways. First, it uses them to rank and select the candidate semantic concepts from the semantic representation of input for mutation. Second, it uses these concepts to direct the mutation in the area of most attention of the DL system, enabling the approach to perform targeted input modifications.
[…]
Our paper makes the following contributions:
Approach.
A mutation operator for semantic-based fuzzing of DL systems based on local explanations that XAI techniques produce.
Technique.
An implementation of our approach for focused fuzzing of DL systems, implemented in the tool XMutant. To the best of our knowledge, this is the first solution that uses XAI techniques for semantic-based focused test generation, both at the model and system level.
Evaluation.
An empirical study on three different DL systems showing that XMutant is better than a state-of-the-art approach, in terms of higher effectiveness, efficiency, validity, and label-preservation rates.
2 Background
Deep Neural Networks (DNNs) are increasingly used in complex safety-critical tasks, such as autonomous driving nvidia-dave2 , autonomous aviation Julian , medical diagnosis zhang2018medical or disease prediction Zhao-nature . This paper leverages XAI techniques for testing DL systems, i.e., systems that use DNNs. [^1] This paper uses the terms DNNs and DL systems interchangeably, for simplicity of exposition. In the following, we provide background on foundational concepts used in this paper: (1) semantic-based input representation and (2) XAI methods.
2.1 Semantic-based Input Representation
Semantic-based test generation approaches utilize abstract aspects of the inputs, such as the shape of a digit in an image or the sentiment polarity of a word in a text, rather than relying on concrete aspects of the inputs, such as changes of pixel values in an image or letter modifications in a word deepatash ; 2021-Riccio-ASE ; zohdinasab2021deephyperion . Semantic-based approaches include model-based techniques 2020-Riccio-FSE , which utilize domain-specific models to generate test inputs. These methods can translate the input from its concrete representation to its abstract representation and back. By manipulating inputs in this abstract space, perturbations gain semantic meaning, allowing for controlled perturbations that reflect meaningful changes in the original data. Semantic-based input representation brings an important benefit for test generation–it reduces the dimensionality of the search space. Conceptually, there exists a 1-to-many mapping from abstract to concrete values, enabling the smaller input space to be more exhaustively covered. Furthermore, semantic validity criteria are expressed at the model level, preventing automated algorithms from creating unrealistic inputs that fall outside the valid operational domain of the DL system 2023-Riccio-ICSE ; jiang2024validity ; zhang2024enhancing .
2.2 XAI Methods
XAI methods are used to make the decisions of complex DL systems more transparent and understandable vilone2020explainable ; gunning2019xai ; dovsilovic2018explainable , which is essential for validation, trust-building, and regulatory compliance. In this paper, we consider XAI methods and use cases (sentiment analysis, digit recognition, and advanced driving assistance systems or ADAS) established in prior work on DL testing 2020-Riccio-FSE ; 2021-Riccio-ASE ; deepatash .
For sentiment analysis, we consider the local explanations produced by LIME lime , SmoothGrad smilkov2017smoothgrad , and Integrated Gradients integratedgradients . In brief, LIME lime , namely local interpretable model-agnostic explanation, explains the predictions of DL systems by approximating their behavior with a simpler, interpretable, surrogate model in a local region. However, a large number of perturbations and inferences are required in this process, which makes LIME computationally expensive. Therefore, we also consider other gradient-based XAI methods (SmoothGrad and Integrated Gradients) to ensure the efficiency of test generation. These approaches are representative of different families of XAI methods and address known issues of saliency methods, such as gradient discontinuity and saturation. Particularly, SmoothGrad smilkov2017smoothgrad reduces the noise of gradient-based explanations by adding artificial noises and averaging them, whereas Integrated Gradients integratedgradients overcomes the gradient saturation by summing over scaled inputs.
Handwritten digit recognition and advanced driving assistance are imagery tasks, for which we consider saliency or pixel attribution methods simonyan2013deep ; shrikumar2017learning ; li2021experimental . These solutions generate local explanations in the input space, identifying regions of an input image that influence the decision-making process of DNNs. Particularly, in addition to using SmoothGrad smilkov2017smoothgrad and Integrated Gradients integratedgradients mentioned before, we also considered CAM (Class Activation Map)-based methods, as the models under test are typically convolutional neural networks. For example, Grad-CAM++ Grad-CAM++_2018 generates heatmaps by computing a weighted combination of the positive partial derivatives from the last convolutional layer.
3 XMutant
Table 1: Testing characterization for three popular tasks often solved with DL systems.
Sentiment analysis Digit recognition Advanced driving assistance
Testing level Model Model System
DNN task Classification Classification Regression
Test input Text Digit image Road in a driving simulator
Input to DL system Word tokens Digit image Driving scenario
Semantic representation Words Control points (image) Control points (road center)
Failure criterion Misclassification Misclassification Out of road bounds
XMutant is a focused fuzzing technique for DL systems. It leverages post-hoc local explanations for individual predictions to guide fuzzing. XMutant is applicable to any DL system where semantic representations are available. As mentioned in Section 2.1, we instantiate XMutant for tasks established in prior work on DL testing 2020-Riccio-FSE ; 2021-Riccio-ASE ; deepatash for which semantic models exist: (1) sentiment analysis, (2) digit recognition, and (3) advanced driving assistance systems (ADAS).
Testing scenarios.  Table 1 shows the characteristics of our testing scenarios. Semantic representations are often domain-specific. For example, both sentiment analysis and digit recognition are classification tasks, but the semantic modelings are different. In sentiment analysis, the words in a movie review are mapped to an embedding space whereas in digit recognition, the bitmap of a MNIST handwritten digit mnist is converted into sequences of cubic Bézier curves farin1983algorithms defined by a series of control points, according to the Scalable Vector Graphics representation. For advanced driving assistance, the centerline of a road in a driving scenario is represented by Catmull-Rom cubic splines catmull1974class , also specified by a sequence of control points. The different testing levels also imply distinct interpretations of failures. For sentiment analysis and digit recognition, a failure is defined as the disagreement between the output of the DNN under test and the ground truth, e.g., a misclassification of the movie review or the handwritten digit. For advanced driving assistance, failures are identified by the misbehavior of the whole DL system, i.e., the autonomous vehicle, in response to the outputs of the DNN, rather than isolated incorrect predictions on individual images. Following existing research Abdessalem-ICSE18 , we evaluate whether a sequence of incorrect DNN predictions leads to a violation of the safety requirements of the system, i.e., the vehicle driving out of the road, or a reduction of the driving quality 2021-Jahangirova-ICST .
Input: DNN under test, set of test inputs $T$, failure criterion, termination condition.
Output: Failure set $F$
1
2Initialize test inputs $T$ with random individuals if $T=\emptyset$;
3 Initialize failure set $F=\emptyset$;
4 while $T\neq\emptyset$ or termination condition do
5       for each test input $t$ in $T$ do
6             Test DL system with the input $t$;
7             if failure criterion not met then
8                   $S$leftarrow GetSemanticRepresentation($t$) ;
9                   $e$leftarrow LocalExplanationComputation($t$) ;
10                   $cs$leftarrow SemanticConceptSelection($S$, $e$) ;
11                   $d$leftarrow MutationDirectionComputation($S$, $e$, $cs$) ;
12                   $S$leftarrow Mutate($S$, $cs$, $d$) ;
13                   $t$leftarrow GenerateConcreteInputFromSemantic($S$) ;
14
15            else
16                   Move $t$ from $T$ to $F$;
17
18             end if
19
20       end for
21
22 end while
23return $F$;
Algorithm 1 XMutant: XAI-guided Fuzzing
Algorithm.  Algorithm 1 shows the main steps of XMutant. XMutant takes as input the DL system, a set of test inputs $T$, the failure criterion, and a termination condition, such as a time budget or a number of generated inputs. XMutant produces as output a set $F$ containing failure-inducing inputs on the DNN under test. The test inputs in $T$ are used to test the DL system. Initially, XMutant initializes the set of test inputs $T$, e.g., by randomly selecting seeds from a test dataset, when they are available (for sentiment analysis or digit recognition), or by randomly generating inputs if they are unavailable (for the advanced driving assistance system). If an original test input $t$ produces a failure, it is discarded. The main loop of the algorithm (Lines 7—12) retrieves the semantic representation $S$ (Line 7), which is a sequence of semantic concepts, and the local explanations $e$ (Line 8), which will be used for focused mutations. Then, a candidate semantic concept $cs\in S$ is selected for mutation (Line 9), based on the area of highest attention indicated by the local explanation $e$, followed by a mutation of the semantic concept in such area (Lines 10—11). Finally, the newly generated semantic-based representation is restored to the original input space (Line 12) and used for testing the DL system. In case of failure, the algorithm stores the evolved test input $t$ to the failure set (Lines 14). Finally, the algorithm returns the failure set $F$.
3.1 XAI-guided Mutation Operator
XMutant’s insight is to fuzz test input on the critical areas of the DNN attention to create challenging inputs to the corresponding DL system and, consequently, trigger misbehaviors. XMutant’s mutation operator selects the semantic concept that is near a high-attention region and, subsequently, performs a targeted mutation in such region.
XMutant’s mutation operator comprises three steps: (1) computing the local explanation (Line 8), (2) selecting the semantic concept to mutate (Line 9), and (3) determining the mutation direction (Line 10). First, XMutant uses an XAI technique to compute local explanations associated with a DNN and its test instances (procedure LocalExplanationComputation on Line 8). Typically, the explanation of the DNN prediction for a given input has the same size as the input, with each dimension of the input space corresponding to an explanation score. Second, XMutant selects the semantic concept to mutate on the high-attention area of the local explanation (procedure SemanticConceptSelection on Line 9), where an intermediate weight vector ${w}$ is obtained and denotes the importance of semantic concepts with $dim({w})=dim(S)$. For textual inputs, the local explanation corresponds directly to the semantic concept, hence selecting the semantic representation is based on the explanation’s magnitude (i.e., ${w}=e$ and $dim(e)=dim({w})$). However, for model- or system-level imagery inputs, our approach needs to map the high-dimensional explanations, onto the lower-dimensional semantic concept space through a function $f$ before the selection, i.e., ${w}=f(e)$ and $dim(e)>dim({w})$. To ensure diversity of generated test inputs, XMutant randomly selects candidate semantic concepts $cs$ based on the weights ${w}$, where semantic concepts with higher weights have higher chances of being chosen Third, when a candidate semantic concept is selected, XMutant leverages the local explanations to choose the mutation direction, aiming to maximize the mutation effectiveness (Line 10).
XMutant is configured according to the kind of input and semantic representation associated with the task under test (Table 1). We elaborate on the domain-specific configurations in terms of semantic concept selection and mutation direction in the following sections.
3.2 Model-level Application on Textual Inputs
For textual inputs, the inputs to the DNN are tokens corresponding to the words. Therefore, obtained explanations have the same dimensions as the semantic representation, i.e., words.
With model-agnostic XAI technique methods such as LIME lime , the feature attributions for each token can be computed directly. With gradient-based XAI methods, however, the presence of a non-differentiable embedding layer in the DL model under test, makes the gradient computation unavailable. To address this, we propose a customized approach for interpreting such DL models. Let us assume that the embedding layer mapping the input vector $x$ consisting of $L$ tokens in the embedding space of dimension $N$, is a matrix of size $\mathbb{R}^{L\times N}$. Then we take the following steps. First, we remove the embedding layer from the model under test to obtain the corresponding submodel. Second, we apply gradient-based XAI methods to compute the explanation $\hat{e}\in\mathbb{R}^{L\times N}$ in the embedding space for the submodel, with each row vector $\hat{e}[i]\in\mathbb{R}^{N}$ representing the explanation vector for the corresponding token $x[i]$. Third, we map the explanation in the embedding space back to the input space with summation (for Integrated Gradients) or norm operators (for SmoothGrad), e.g., $e[i]=sum(\hat{e}[i])$ for each token $x[i]$. Finally, we obtain a local explanation $e$ that has the same dimension as the input $x$.
The local explanations obtained for textual inputs differ in their characteristics depending on the XAI method used. For instance, SmoothGrad captures the gradient information, indicating the sensitivity of the predictions to word variations (Figure 1 (a)). In this case, the weight vector consists of absolute gradient information. In contrast, LIME and Integrated Gradients provide explanations on feature attributions, which represent the magnitude of the positive (or negative) contribution of a word on the prediction, making them useful for guiding the mutation direction (Figure 1 (b)). In this case, the weight vector contains the raw explanations.
Figure 1: Example of movie review sentiment analysis case study (best viewed in color). (a) Local explanation by SmoothGrad; (b) Local explanation by LIME.
Regarding the actual mutation operation, we used two mutation methods available in the literature deepatash : (1) replacing a word with its synonym obtained from WordNet miller1995wordnet , and (2) adding an “and” conjunction after an adjective (or adverb), followed by a synonym of the adjective (or adverb). These mutation methods ensure that the original meaning of the sentence is preserved, thereby maintaining a high validity of generated test inputs.
During the fuzzing process, one of the mutation methods is selected at each iteration. When the first mutation method is chosen, we apply it to the selected word, regardless of the considered XAI method. This is because it is uncertain whether replacing a synonym will increase or decrease the contribution of the selected word. In contrast, the second mutation method enhances the semantics of the sentence by adding a synonym. Therefore, for XAI methods that differentiate feature attributions (i.e., LIME lime and Integrated Gradients integratedgradients ), we perform targeted mutation by applying the second mutation method only to words with opposite effects to the prediction (e.g., we modify “promptly” to “promptly and quickly” but we do not modify “overwhelmed” in Figure 1), which helps to more effectively challenge the model under test. For XAI methods that are not focused on explaining the feature attributions (i.e., SmoothGrad), we apply the second mutation method on the selected word, regardless of its effect on the prediction.
3.3 Model-level Application on Imagery Inputs
For imagery inputs, the local explanation, so-called heatmap, is itself an image with the same dimensions as the original image, where pixels are highlighted based on the attention that the DNN under test gives to each portion of the image. As such, both the selection of the candidate semantic concept to be mutated and the computation of the mutation direction are based on a single heatmap $\mathbf{H}$.
Figure 2 shows the different steps of the XMutant operator for a handwritten MNIST mnist digit image. Given an image, XMutant converts the bitmap representation into an SVG representation to extract the semantic representation, i.e., a sequence of control points (Figure 2 (a)). Then, it computes the heatmap on this SVG representation using an XAI method (e.g., Grad-CAM++ in Figure 2 (b)); in the heatmap relevant locations correspond to hot color intensities (e.g., red/yellow), whereas irrelevant locations correspond to cold color intensities (e.g., blue). The heatmap values are normalized such that low-attention values are close to zero, while high-attention values are close to one.
Selecting the candidate semantic concept, i.e., the control point $cp_{k}$, for mutation requires the computation of the weights corresponding to the control points based on the heatmap. We evaluate two techniques in this paper, based on square windows and clustering.
3.3.1 Square Windows Semantic Concept Selection
Square windows involve computing the weight of a control point by averaging the intensity values within a square window centered on the control point $(x_{k},y_{k})$ (e.g., the red squares in Figure 2 (c)), according to the following equation:
$w_{k}=\frac{1}{(2d+1)^{2}}\sum_{i=-d}^{i=d}\sum_{j=-d}^{i=d}\mathbf{H}[(\mathbf{p}_{c_{k}}+(i,j))]$ (1)
where $d$ is the distance between the center of the square $(x_{k},y_{k})$ and each side (and the window size ws is given by $\textit{ws}=2d+1$), while $\mathbf{p}_{c_{k}}\in\mathbb{R}^{2}$ denotes the coordinates of the control point in the heatmap. Subsequently, the weights ${w}$ for each semantic concept are normalized.
The square windows approach, although straightforward, involves choosing an appropriate window size ws. When the window size is large, it can exceed the input’s boundaries or have squares of different control points overlap. If the window size is small, large regions of the local explanation might be uncovered, potentially missing some high-attention areas. Additionally, this method does not provide any guidance for determining the direction of mutation.
Figure 2: XMutant steps applied to the MNIST case study (best viewed in color). (a) SVG semantic-based representation of a digit 3. Control points (i.e., semantic representation) are shown as red crosses; (b) Heatmap generated by Grad-CAM++; (c) Square windows centered on the control points; (d) Clustering on the local explanation (one cluster per control point); (e) Control point selection (red cross marker) and mutation towards the attention attractor of the cluster (green star); (f) Digit 3 resulting by mutating the candidate control point.
3.3.2 Clustering-based Semantic Concept Selection
To overcome the limitations of the square windows approach, we also propose a control point selection technique based on clustering, which is less reliant on hyperparameters and, thus, more automated.
XMutant uses the k-means clustering algorithm wu2012advances to cluster pixels based on their values with the number of clusters $n$ to match the number of control points. We modified the typical k-means process by initializing the centroids with the control points and limited the algorithm to a single iteration, without convergence. Figure 2 (d) shows 13 clusters identified by different colors, one for each control point. Then, XMutant computes the weight $w_{k}$ for each control point $cp_{k}$ as the sum of the ratios between each pixel value in the cluster of $cp_{k}$, and the distance between the coordinates of the pixel and the respective control point:
$w_{k}=\sum_{i=1}^{m_{k}}\frac{\mathbf{H}[\mathbf{p}_{i}^{k}]}{d_{i}^{k}},\text{ with }d_{i}^{k}=\max(1,||\mathbf{p}_{i}^{k}-\mathbf{p}_{cp_{k}}||)$ (2)
where $m_{k}$ is the number of pixels in the cluster of control point $cp_{k}$, $\mathbf{p}_{i}^{k}\in\mathbb{R}^{2}$ denotes the coordinates of $i$-th pixel in the cluster, and $d_{i}^{k}$ is the Euclidean distance between $\mathbf{p}_{i}^{k}$ and the coordinates of the control point in the heatmap (i.e., $\mathbf{p}_{cp_{k}}$). To avoid division by zero when $\mathbf{p}_{i}^{k}=\mathbf{p}_{cp_{k}}$, we set the lower bound of $d_{i}^{k}$ to one.
Once the weights have been normalized and the control point $cp_{k}$ sampled (e.g., the control pointing the green cluster in the bottom-left corner of Figure 2 (e)), XMutant mutates the selected control point towards the attention attractor $\mathbf{c}_{k}$, which is defined as the center of intensity of the respective cluster, given by:
$\mathbf{c}_{k}=\frac{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]\mathbf{p}_{i}^{k}}{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]}$ (3)
Figure 2 (e) shows an example of how the selected control point in the green cluster (bottom left) is mutated towards the respective attention attractor. This control point is marked by a red cross and moves toward the green star, which represents the attention attractor of the green cluster. The mutation direction is depicted by a black arrow extending from the control point to the attention attractor, leading to a contraction of the high-attention area. Figure 2 (f) shows the resulting digit after the mutation, where the control point in the bottom-left corner of the image (see Figure 2 (a)) has shifted further downward.
3.4 System-level Application on Logical ADAS Scenarios
As the ADAS processes a stream of images from a driving simulator, applying our approach at an individual image level (as described previously for model-level testing) does not effectively assist in selecting and mutating candidate semantic concepts, i.e., control points. Instead, XMutant generates a series of heatmaps for each image perceived by the DNN, and aggregates them to retrieve a semantic score that reflects the system’s overall performance.
Figure 3 shows the steps of XMutant for the ADAS. The test input is a sequence of control points located at the center of the road (Figure 3 (a)), determining the shape of the two-lane road that will be instantiated in the driving simulator (we assume the other environment variables, such as weather conditions, as fixed). Once the road is instantiated, the DNN that controls the vehicle takes as input the frames that are recorded by the onboard camera (Figure 3 (b)) and outputs the driving commands (i.e., steering angles). Therefore, XMutant computes the heatmaps of such camera frames, resulting in a sequence of heatmaps (Figure 3 (c)).
Figure 3: XMutant steps for the ADAS (best viewed in color). (a) A road with 3 control points, or semantic representation (red crosses). The road sectors are indicated as Roman numerals and separated by dashed cyan lines; (b) A sequence of driving frames recorded by the vehicle’s camera; (c) A sequence of heatmaps, corresponding to the driving frames, generated by SmoothGrad; (d) Weights for each road sector and associated semantic concept; (e) Control point selection (top-right) and mutation towards the right lane (dashed yellow line indicates the new centerline); (f) Road resulting from the mutation of the control point in sector III.
To weigh the control points, we divide the road into segments extending from each control point to the road’s centerline (e.g., in Figure 3 (a) there are three road sections). Each road segment is initially linked to the control point at its start. This approach is based on preliminary findings indicating that mutations at a control point predominantly affect the driving behavior in the section immediately following it. To aggregate sequences of heatmaps, XMutant computes the derivative of a heatmap $\nabla\mathbf{H}_{t}$ at a certain time step $t$ as the average absolute difference between the pixels of two consecutive heatmaps $\mathbf{H}_{t-1}$ and $\mathbf{H}_{t}$:
$\nabla\mathbf{H}_{t}=\frac{1}{wh}\sum_{i=1}^{w}\sum_{j=1}^{h}|\mathbf{H}_{t-1}[i,j]-\mathbf{H}_{t}[i,j]|$ (4)
Subsequently, the sampling weight $w_{k}$ of the control point $cp_{k}$ is given by the mean value of all derivatives in the corresponding road section, i.e., $w_{k}=1/m_{k}\sum_{i=2}^{m_{k}}\nabla\mathbf{H}_{i}^{k}$, where $m_{k}$ is the number of frames in a specific road section, and $\{\mathbf{H}_{1}^{k},\dots,\mathbf{H}_{m_{k}}^{k}\}$ denotes the heatmaps in the road section associated to the control point $cp_{k}$.
Indeed, existing work 2022-Stocco-ASE has shown that the derivative of a local explanation is indicative of a poor-performing DNN. In this way, XMutant grants a higher weight to road sections—and associated control points—where the driving behavior of the DNN is more erratic. Consequently, by mutating such control points in the corresponding road sections, XMutant is more likely to generate a test input that induces more challenging driving conditions. For instance, in Figure 3 (d), the road section with the highest mean derivative is road section III; in this example, the corresponding control point in the top-right corner of Figure 3 (e) is sampled.
Ultimately, XMutant determines the mutation direction by considering the perpendicular line to the road’s centerline at the candidate control point. Consequently, only two directions are possible: moving the control point toward the right lane or the left lane.
To determine the appropriate mutation direction, we analyze the series of heatmaps for the chosen road segment. XMutant partitions each local explanation into two equally-sized sub-heatmaps, i.e., $\mathbf{H}[:w/2,:]$ and $\mathbf{H}[w/2+1:,:]$ (where the column separates start and end indices for each dimension of the local explanation). XMutant then calculates the difference in average intensity between the left and right sub-heatmaps in the road segment. The mutation direction is chosen based on the side with the highest average intensity (in Figure 3 (e) the selected direction is the right direction, hence the last control point is translated towards the right lane, resulting in the test input shown in Figure 3 (f)).
4 Empirical Study
4.1 Research Questions
We consider the following research questions:
RQ1 (effectiveness): How effective is XMutant at finding misbehaviors?
RQ2 (efficiency): How efficient is XMutant at finding misbehaviors?
RQ3 (configuration): How do effectiveness and efficiency vary when considering different XAI algorithms? What is the best configuration of semantic concept selection and mutation direction?
RQ4 (validity): To what extent are the inputs generated by XMutant valid and label-preserving?
RQ5 (comparison): How does XMutant compare with gradient-guided raw input manipulation techniques?
The first research question (RQ1) assesses whether XMutant attains a high failure rate. The second research question (RQ2) evaluates the efficiency of XMutant. The third research question (RQ3) evaluates different configurations of XMutant obtained by varying the XAI algorithm, semantic concept selection, and mutation direction strategy. The fourth research question (RQ4) studies the usefulness of the inputs that XMutant produces, both for automated input validators and human assessors. The final research question (RQ5) examines how XMutant compares to widely used gradient-guided raw input manipulation techniques. Although these techniques differ in their primary objectives, generalization and robustness testing are often confused since they share the same goal, i.e., DNN reliability. Thus, we analyze and compare the generated inputs from our approach and those generated by existing methods to provide evidence of the benefits of a semantic-based representation, as test inputs generated by these distinct methodologies may form different manifolds in the semantic space.
4.2 Objects of Study
4.2.1 Datasets and Models
IMDB. Concerning sentiment analysis, we consider a DL system designed to classify the sentiment (i.e., positive or negative) of movie reviews from the IMDB database imbd . We use the DL system available in the replication package of existing work man_sentimental_imdb_lstm , characterized by an embedding layer, an LSTM layer hochreiter1997long , and two fully connected layers. The model achieves 85.25% accuracy on the test set.
MNIST. Concerning digit recognition, the DL system classifies handwritten digits from the MNIST dataset lecun1998gradient . This DL system takes 28x28 greyscale images as input and predicts the corresponding digit (the possible classes range from 0 to 9). In this paper, we test the convolutional DNN architecture provided in the replication package of existing work 2020-Riccio-FSE . Architecturally, it is characterized by 2 convolutional layer, a max-pooling layer, and two fully connected layers. The model achieves 98.99% accuracy on the test set.
ADAS. Concerning advanced driving assistance, the DL system controls a vehicle in the Udacity simulator udacity-simulator , a cross-platform driving simulator developed with Unity3D unity , widely used in the ADS testing literature 2020-Stocco-ICSE ; 2020-Riccio-EMSE ; 2021-Jahangirova-ICST ; 10.1007/s10515-021-$0-$ The DL system includes a DAVE-2 model (a DNN regressor), a Level 2 iso26262 ADAS that performs the lane keeping functionality from a training set of images collected when the driver is an expert human pilot, by predicting the corresponding driving commands imitating the human driving behavior. We obtained the trained DAVE-2 model and the simulator from the replication package of existing work 2024-Biagiola-EMSE . The model architecture includes three convolutional layers for feature extraction, followed by five fully connected layers. The simulator supports the creation of open-loop road tracks for testing ADAS models, including the ability to generate and modify road topologies.
4.2.2 Comparison Baseline
To assess the guidance provided by our XAI-based approach, we compare XMutant against DeepJanus, a popular semantic-based test generator 2020-Riccio-FSE . For the selection of semantic concepts and mutation directions, DeepJanus relies on a random method that selects a candidate semantic concept for mutation based on a uniformly weighted probability. The mutation direction is task-specific. For IMDB, it applies one of the mutation methods on a randomly selected semantic concept. For MNIST, the mutation direction is sampled from the uniform distribution over 00 to $2\pi$. For ADAS, the mutation direction is randomly selected from one of the two directions perpendicular to the road’s curvature. To ensure a fair comparison, we applied the same extent value for mutation direction to both XMutant and DeepJanus’s mutation methods. In our study, we sampled the extent of MNIST from a uniform distribution over the interval $[0,1.2]$, and for ADAS we adopted a fixed value 4, proportional to the road’s width 8. These values are determined based on existing literature 2020-Riccio-FSE ; 2024-Biagiola-EMSE .
Regarding raw input manipulation approaches, we selected popular approaches that use gradients to guide the generation of adversarial inputs, i.e., FGSM Goodfellow2014ExplainingAH , DeepXplore pei2017deepxplore , and DLFuzz guo2018dlfuzz . For FGSM, we progressively increased the perturbation intensity up to the 20% threshold, as higher intensities affect the validity of the generated inputs (i.e., the original images were no longer recognizable due to excessive noise or corruption). For DLFuzz and DeepXplore, we adopted the default settings described in the original papers.
4.2.3 Metrics used for Analysis
Concerning RQ1, we evaluate the effectiveness of the test generation technique by computing the cumulative failure rate observed under a given number of mutation iterations across different configurations. The cumulative failure rate is obtained by dividing the number of failures by the total number of inputs. For IMDB and MNIST, a failure is characterized by the number of misclassifications. For ADAS, we measure the number of safety-critical failures in terms of out of bounds (OOBs), which occur when the vehicle drives outside the road’s drivable lanes during the execution in the simulator.
Concerning RQ2, we evaluate the performance of XMutant by measuring the relative efficiency with respect to DeepJanus. Particularly, relative efficiency is computed by integrating the cumulative failure rates over the number of iterations to obtain the area under the cumulative failure curve (AUFC). For each configuration and number of iterations, we calculate the AUFC (Area Under the Failure Curve) and then divide it by the AUFC of our comparison baseline. Considering computational time complexity, XMutant introduces an additional overhead for computing local explanations compared to the baseline. To assess its impact, we specifically evaluated the proportion of the XAI overhead relative to the baseline’s average computation time per iteration.
Concerning RQ3, we assess the effectiveness and efficiency of each configuration of XMutant using the metrics used for RQ1 and RQ2. In RQ4, we present the validity rates for both automated input validation and human evaluation, along with the label-preservation rates as judged by human assessors.
In RQ5, we evaluate efficiency and effectiveness of competing approaches, by calculating the number of misclassified inputs divided by the total number of original seeds, and the total elapsed time divided by the number of misclassified inputs. We also adopt two metrics to evaluate the realism of generated inputs. The density and coverage metrics ferjad2020icml measure how well the generated inputs align with real data distribution, using k-nearest neighbors to define local manifolds. Particularly, coverage quantifies the proportion of real points with nearby generated points, while density reflects the concentration of generated points around the original test inputs. Following the reference implementation in Dola et al. dola2024cit4dnn we set $k=5$.
4.2.4 Configurations
We evaluate a total of 24 configurations, varying the XAI algorithm, the semantic concept selection method, and the mutation direction as described next.
For IMDB, XMutant uses three XAI algorithms (SmoothGrad smilkov2017smoothgrad , LIME lime , and Integrated Gradients integratedgradients ) and two mutation methods from the literature, namely synonym replacement and addition zohdinasab2021deephyperion .
In the MNIST classification task, XMutant uses three XAI algorithms (SmoothGrad smilkov2017smoothgrad , Grad-CAM++ Grad-CAM++_2018 , and Integrated Gradients integratedgradients ) and two strategies (square windows and clustering). For square windows, we use a window size value of $\textit{ws}=3$, which was found appropriate during preliminary experiments (e.g., such value does not lead to the creation of square windows that exceed the input boundaries). For mutation direction, XMutant relies on random selection due to the absence of attractor information. The clustering strategy allows for the mutation direction to be chosen either towards the attention attractor or randomly. Additionally, XMutant pre-processes the heatmap to discard low-intensity values and filter out the noise using a threshold $\epsilon=0.1$, which was found adequate from preliminary experiments.
For the ADAS regression task, XMutant uses three XAI algorithms (SmoothGrad smilkov2017smoothgrad , Grad-CAM++ Grad-CAM++_2018 , and Integrated Gradients integratedgradients ). Following existing literature, it adopts the heatmap derivative function method from ThirdEye 2022-Stocco-ASE to aggregate consecutive heatmaps into a single score. Considering the mutation direction, XMutant uses three strategies (the attractor mutation direction does not apply to ADAS). The first strategy determines mutation directions by identifying the lane receiving the most focus (High). For instance, if the heatmaps indicate higher attention on the right lane, the mutation direction chosen is to the right. In contrast, the second strategy involves mutation directions in lanes receiving the least focus (Low). For example, if the heatmaps reveal the right lane as receiving the most focus, the mutation direction is set to the left. The final strategy adopts a random approach to select the mutation direction (i.e., either left or right).
4.2.5 Procedure
Concerning RQ1, RQ2, and RQ3, the evaluation procedure is as follows.
For IMDB, we randomly select $200$ movie reviews from the original IMDB test set. We discard $26$ reviews that are misclassified before mutation. For MNIST, we randomly select 2,000 digits from the original MNIST test set, uniformly distributed for each class. We run the classifier to ensure that all original seeds are correctly classified. We discard $23$ digits identified during this screening task and evaluate the remaining seeds. For ADAS, we randomly generate $60$ test inputs (i.e., roads) with the following characteristics: maximum curvature $70$ degrees and $12$ control points to avoid too many invalid roads and ensure the variety of choice during mutation 2024-Biagiola-EMSE . As a sanity check, we ensured that the trained DAVE-2 model drives all roads in the simulator successfully. As a result, we discard $5$ roads that trigger misbehaviors before any mutations. We execute each configuration of XMutant and DeepJanus, using a budget of $100$ iterations for IMDB, 1,000 for MNIST, and $30$ for the ADAS. This upper bound value was found through experimentation to be adequate for convergence of either XMutant or DeepJanus.
Concerning RQ4, we assess the validity of XMutant’s output using automated validators and questionnaires for human assessment. Particularly, for IMDB, we perform the validity check using ChatGPT openai , a state-of-the-art large language model. For textual inputs, manual validation is inherently subjective and can lack consistency. Additionally, the workload for manual evaluation is significant, potentially leading to decreased evaluation quality. Large language models like ChatGPT, on the other hand, provide a viable automated solution for validation and are increasingly regarded as reliable as humans for comprehending and explaining small chunks of text, such as those produced by XMutant for IMDB, as highlighted by recent studies wang2025llmsreplacehumanevaluators ; zheng2023judgingllmasajudgemtbenchchatbot .
We used ChatGPT-4o-mini GPT since it is cost-efficient and shows strong performance on textual intelligence and reasoning tasks. We append the textual input (to be classified) to the following prompt “Assume you are a sentiment classifier, given a text of a movie review removing Stopwords and Punctuations. Please only reply ‘positive’, ‘negative’, or ‘invalid’ if the sentence does not make sense.”. Due to the inherent randomness of ChatGPT, we repeated the validity assessment five times for each input and reported the average score.
For MNIST, we randomly select 200 misclassified inputs produced by the best configuration of XMutant and the DeepJanus baseline. We assess the validity of XMutant’s and DeepJanus’s output using SelfOracle 2020-Stocco-ICSE , a distribution-aware input validator for imagery data, which has shown high agreement with human validity assessment in a large comparative study about test input generators for DL 2023-Riccio-ICSE . We obtained the trained model of SelfOracle for MNIST from the replication package of the paper by Riccio and Tonella 2023-Riccio-ICSE . We applied SelfOracle to reconstruct all failure-inducing images by each variant of XMutant and DeepJanus using the same rate of false alarms as the original study ($\epsilon=0.05$%). We also evaluate validity and label preservation with human assessors. We conducted a questionnaire where 10 participants were asked to identify a digit, with choices ranging from 0 to 9, or to indicate if it was unrecognizable as a digit. This method allowed us to assess both the validity (if the response is a digit) and label preservation (if the response matches the intended label).
For ADAS, we did not perform any automated or manual validation of the roads, as our test generation process filters out invalid roads by ensuring they meet specific criteria: (1) the start and end points are different; (2) the road is contained within a square bounding box of a predefined size (specifically 250 $\times$ 250), and (3) there are no intersections.
Concerning RQ5, we conduct the comparison only for the image classification task, as the comparing techniques do not apply to textual inputs and logical driving scenarios. We use the same original 2,000 seeds for all techniques, for which we compute the effectiveness and efficiency metrics. For coverage analysis, we randomly selected 50 generated failure-inducing inputs per class, totaling 500, to ensure that coverage-related metrics are not affected by various population sizes. However, for DeepXplore, we only obtain 44 generated inputs out of 2,000 original seeds, so we analyze all available samples.
Since our analysis prioritizes semantic similarity instead of pixel-wise similarity, we conduct coverage analysis in an embedding space that captures high-level semantic features. We use a pretrained VGG16 model on ImageNet as an embedding extractor, as it has demonstrated 100% accuracy in MNIST classification through transfer learning kaggle_vgg16_mnist , confirming its ability to capture relevant digits features. The 512-dimensional embedding vectors are then visualized by reducing their dimensionality using principal component analysis (PCA), to show the overlap between generated and real data distributions.
4.3 Results
4.3.1 Effectiveness, efficiency, configuration (RQ1, RQ2 and RQ3)
Table 2 presents the effectiveness, performance, and comparative results for all configurations of XMutant and DeepJanus as the baseline, for all case studies (IMDB, MNIST, and ADAS). The table reports the information at different iteration intervals, respectively 10 for IMDB, 100 for MNIST, and 5 for ADAS. For each configuration of XMutant and DeepJanus, the first number denotes the cumulative failure rate, whereas the second number indicates the relative efficiency with respect to the baseline.
Concerning effectiveness (RQ1), for all case studies, all configurations of XMutant outperform the baseline, regardless of the iteration considered. We assessed the statistical significance of the differences in the cumulative failure rate between XMutant and the baseline using the non-parametric Mann-Whitney U test Wilcoxon1945 (with $\alpha=0.01$) and the magnitude of the differences using the Cohen’s $d$ effect size cohen1988statistical . The differences were found to be statistically significant ($p$-value $<$ 0.01), with varying effect sizes.
Table 2: Results for RQ1, RQ2 and RQ3. For each configuration, the two values indicate the Cumulative Failure Rate [%]/ Relative Efficiency over different Iterations. P-values indicating statistically significant differences are highlighted in bold.
Iterations SmoothGrad LIME Integrated Gradients DJ
10 14/1.93 11/1.44 16/2.05 7/-
20 22/2.08 32/2.31 31/2.71 10/-
30 31/2.08 42/2.74 46/2.92 14/-
40 39/2.12 51/2.82 57/3.05 18/-
50 41/2.12 60/2.86 64/3.09 21/-
60 47/2.08 63/2.83 67/3.04 24/-
70 49/2.03 66/2.75 70/2.94 27/-
80 51/1.98 70/2.69 74/2.87 29/-
90 53/1.93 72/2.64 76/2.80 31/-
IMDB 100 56/1.90 76/2.74 76/2.59 32/-
p-value 6.93E-18 1.17E-17 7.90E-18 -
RQ1 effect size large large large -
p-value 2.53e-04 1.20e-06 2.92e-06 -
RQ2 effect size medium large large -
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Clustering SW Clustering SW Clustering SW -
Iterations Attractor Random Random Attractor Random Random Attractor Random Random -
100 44/7.00 11/1.83 11/2.11 36/4.81 07/1.03 09/1.32 34/5.54 13/2.06 13/2.15 06/-
200 77/6.37 24/1.77 24/1.82 74/5.54 21/1.32 22/1.48 60/4.92 27/2.00 26/1.97 15/-
300 88/4.97 38/1.62 36/1.60 88/4.61 35/1.36 34/1.41 74/3.92 41/1.80 37/1.72 26/-
400 94/4.03 47/1.52 45/1.47 94/3.84 46/1.34 45/1.36 80/3.26 51/1.67 47/1.57 35/-
500 96/3.42 55/1.44 53/1.39 97/3.30 54/1.32 54/1.32 85/2.82 59/1.57 54/1.47 43/-
600 97/2.99 62/1.38 60/1.33 98/2.91 60/1.28 61/1.29 87/2.51 66/1.49 61/1.39 51/-
700 98/2.67 67/1.32 65/1.28 98/2.61 66/1.25 67/1.26 89/2.27 71/1.43 66/1.33 58/-
800 98/2.43 71/1.28 69/1.25 99/2.39 70/1.22 73/1.24 90/2.10 76/1.38 71/1.29 62/-
900 99/2.26 75/1.25 73/1.22 99/2.23 74/1.20 77/1.22 91/1.96 79/1.35 74/1.26 67/-
MNIST 1000 99/2.12 78/1.23 76/1.20 99/2.09 77/1.19 80/1.21 92/1.86 82/1.32 77/1.23 70/-
p-value 4.94E-165 1.09E-164 3.46E-165 2.57E-164 1.19E-161 2.10E-164 4.86E-165 7.03E-165 7.05E-165 -
RQ1 effect size large small small large small small large medium small -
p-value 1.75E-183 2.47E-15 1.59E-10 4.46E-188 1.55E-10 2.55E-08 2.06E-142 2.05E-25 3.91E-15 -
RQ2 effect size large small small large small negligible large small small -
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Iterations Random Low High Random Low High Random Low High -
5 27/2.00 31/2.20 42/2.87 49/2.80 42/2.53 36/2.07 45/2.13 36/2.53 40/2.60 22/-
10 65/1.35 60/1.36 67/1.69 75/1.86 87/1.78 75/1.60 64/1.57 65/1.60 58/1.56 51/-
15 84/1.33 76/1.25 82/1.45 85/1.58 93/1.65 95/1.54 75/1.38 78/1.37 80/1.38 65/-
20 89/1.31 82/1.21 84/1.36 95/1.47 93/1.53 95/1.48 87/1.30 85/1.30 85/1.32 75/-
25 91/1.25 85/1.16 93/1.29 95/1.39 95/1.42 95/1.39 91/1.25 87/1.23 89/1.25 84/-
ADAS 30 91/1.21 89/1.13 95/1.24 95/1.32 95/1.35 95/1.32 95/1.21 87/1.18 89/1.20 84/-
p-value 8.23E-08 5.26E-06 1.05E-07 1.07E-07 1.10E-07 1.09E-07 1.05E-07 1.54E-07 1.07E-07 -
RQ1 effect size small small small medium medium medium small small small -
p-value 9.62E-03 2.87E-01 6.93E-03 1.77E-04 1.33E-04 2.19E-04 6.05E-02 5.33E-02 1.47E-02 -
RQ2 effect size small - small medium medium medium - - - -
RQ1: The guidance provided by the XAI allows XMutant to generate significantly more failure-inducing inputs than DeepJanus (up to $+208\%$ for sentiment analysis, up to $+125\%$ for digit recognition, and $+27\%$ for system-level advanced driving assistance).
Concerning performance (RQ2), all configurations of XMutant were producing misbehaviors faster than the baseline, as evidenced by the relative efficiencies regarding DeepJanus as the relative efficiencies are greater than 1. The results hold for all case studies. We also perform the statistical tests for RQ2 on the iteration counts between XMutant and the baseline, which revealed that in 17 out of 21 configurations of XMutant, the differences are statistically significant ($p$-value $<$ 0.01) compared to DeepJanus.
Figure 4 illustrates the time required to compute the local explanations as a percentage of the baseline’s average computation time per iteration. Notably, LIME is extremely time-consuming since it requires estimating a surrogate model, making it unsuitable for test generation. Aside from LIME, XMutant with other XAI methods remains more efficient than DeepJanus, even when the overhead is accounted for. Besides, in system-level ADAS testing, XAI computations constitute a relatively small portion of the total execution time, as executing the driving simulation is computationally expensive. These results show XMutant’s time-efficiency also for complex system-level testing settings.
RQ2: XMutant is faster than DeepJanus at exposing failure-inducing inputs for DL systems (ranging from 2$\times$ to 7$\times$ times faster for model-level sentiment analysis and digit recognition, and 2$\times$ faster for system-level advanced driving assistance).
Figure 4: Computational overhead for local explanation.
Concerning the configurations of XMutant (RQ3), in the case of IMDB, when XMutant is equipped with XAI methods that explain the feature attribution (Integrated Gradients and LIME), it outperforms the baseline both in terms of cumulative failure rate and relative efficiency. However, since the computation time by LIME has shown to be computationally expensive, the configuration using Integrated Gradients is more practical for real-world scenarios.
For MNIST, the performance of XMutant shows negligible differences in both cumulative failure rate and relative efficiency, regardless of the XAI method used, when relying on different control point selection strategies such as Clustering or Square Windows. These configurations consistently achieved a cumulative failure rate over $75\%$ after 1,000 iterations. Notably, when equipped with clustering-based control point selection and random mutation direction, Integrated Gradients proved to be the best configuration. However, the introduction of a mutation direction towards the attractors of the heatmaps significantly enhances XMutant effectiveness. All XAI methods, aside from Integrated Gradients, reached a failure rate of 70% in less than 200 iterations. This performance is on par with that achieved by DeepJanus over 1,000 iterations.
Concerning the ADAS task, the relative efficiency of various configurations with XMutant is approximately twice that of the baseline. However, combining with heatmap-guided mutation direction did not significantly enhance XMutant. Among four different XAI methods, GradCAM++ performed best, achieving a $95\%$ cumulative failure rate in only 15 iterations.
RQ3: For model-level digit recognition, the best configurations of XMutant are SmoothGrad, and Grad-CAM++ equipped with clustering selection and mutation towards the Attractor. For system-level advanced driving assistance, the best configurations of XMutant are Grad-CAM++ with the Random mutation or towards the High attention lane.
4.3.2 Validity (RQ4)
Figure 5 reports the validity rate for IMDB and MNIST. Figure 5 (a) presents the validity assessment by ChatGPT, including the mean and variance due to ChatGPT’s inherent randomness. All configurations of XMutant and DeepJanus maintain a high validity (above 89%) and preservation rate (above 75%) with relatively low deviations since the adopted mutation methods are considered conservative and can preserve the input semantics. DeepJanus has a slightly higher preservation rate as it was assessed on a smaller set of inputs (Section 4.3.1).
(a) IMDB.
(b) MNIST.
Figure 5: Validity and preservation rates for sentiment analysis (left) and digit recognition (right).
For the automated validator of MNIST, we only report the best configurations found in the previous research questions, namely SmoothGrad Clustering Attractor and Grad-CAM++ Clustering Attractor. The plot (Figure 5 (b) top) shows that the validity rate for the DeepJanus steeply decreases with the increasing number of iterations. This is expected because DeepJanus produces out-of-distribution inputs if random mutations are applied an excessive number of times (particularly after 200 iterations). Differently, all configurations of XMutant produce failure-inducing inputs that are regarded as in-distribution for SelfOracle. It is important to notice that the validity rate for XMutant stabilizes at $95\%$ (the used validity threshold in SelfOracle), which indicates that the distribution of the misclassified inputs generated by XMutant highly overlaps with the distribution of the original testing dataset of MNIST. This is attributed to the XMutant’s ability to generate failure-inducing inputs by modifying the original digits slightly, within minimal targeted modifications due to the XAI guidance.
Figure 5 (b) (bottom) reports, at varying iteration steps, the validity and label preservation rates by the human assessors. The plot shows the average value and the variance between assessors. The results from the human assessment differ from those of the automated detector. XMutant still outperforms DeepJanus, with its validity and label-preserving rate converging around $0.9$ and $0.7,$ respectively, after $300$ iterations. In comparison, DeepJanus’s validity and label preservation rate are lower than that of XMutant and exhibit a downward trend.
RQ4: The failure-inducing inputs by XMutant are regarded as valid in-distribution inputs, according to state-of-the-art automated input validators. Moreover, they exhibit a high validity rate ($\approx$90%) and label preservation rate ($\approx$70%), according to human assessors.
4.3.3 RQ5 (Comparison)
Table 3 shows, for the various techniques, the results for effectiveness, efficiency, density, and coverage. In this analysis, we studied the best performing configuration of XMutant (Grad-CAM++ in combination with Clustering and Attractor). XMutant outperforms all other gradient-guided methods also in terms of efficiency and effectiveness. DeepXplore generates a low amount of failures since it only inserts noise in the upper-left occlusion window of the image.
Figure 6 shows that adversarial techniques exhibit a significant shift in the second principal component (PC2) in the PCA space, while the first principal component (PC1) remains relatively aligned with the original data. The failure-inducing inputs generated by XMutant generally adhere to the original distribution and fill in sparsely populated regions of the original distribution (towards the right part). PC1 captures global features such as overall brightness and coarse-grained structure, whereas PC2 is more sensitive to local features such as texture details. Despite minor pixel modifications, RIM techniques seem to generate radically different images in the PCA space. It can be observed that the density and coverage are generally consistent with the PCA visualization, indicating that the inputs generated by the XMutant are closest to the original distribution.
Table 3: Comparison of XMutant and gradient-guided adversarial testing methods.
Technique Effectiveness Efficiency (s) Density Coverage
DLFuzz 60.10% 0.65 0.2156 0.336
FGSM 43.80% 0.37 0.0104 0.05
DeepXplore 0.02% 76.84 0.2591 0.11
XMutant 74.15% 0.33 0.3824 0.41
Figure 6: Semantic space visualization of generated inputs by different techniques.
RQ5: XMutant outperformed 3 gradient-guided adversarial testing methods on four metrics. The failure-inducing inputs by XMutant are able to better cover the original distribution in semantic space.
4.4 Threats to Validity
4.4.1 Internal validity
We conducted comparisons between all variants of XMutant and the baseline within the same experimental framework, using identical benchmarks. A potential internal validity is our implementation of the scripts used to assess these results, which have undergone extensive testing. Additionally, regarding the ADAS model and the simulation platform, we utilized artifacts available from the replication packages of prior studies 2024-Biagiola-EMSE .
4.4.2 External validity
The limited number of DL systems in our evaluation poses a threat in terms of the generalizability of our results. To mitigate this, we considered DL systems addressing different tasks, input data, and testing levels. For the local explanation methods, we considered a limited number of XAI methods. To address this threat, we selected XAI algorithms from diverse families (e.g., surrogate, gradient, and saliency-based methods). Our results show that XMutant consistently outperforms the baseline regardless of the XAI method used. However, it is important to note that the optimal choice of the XAI algorithm is domain-specific.
5 Qualitative Analysis
We analyzed some of the failures qualitatively to understand the reasons why the local explanations are able to guide the mutation process effectively. We focused our analysis on the digit recognition and advanced driving assistance systems.
(a) Qualitative analysis for MNIST by SmoothGrad
(b) Qualitative analysis for MNIST Grad-CAM++
5.1 Qualitative Analysis of Digit Recognition
For model-level digit recognition, we investigate the underlying relationship between model confidence and heatmaps by observing how they change with different mutation configurations. Since heatmaps are high-dimensional matrices, we employ two metrics to extract relevant information. The first metric is the average intensity, for which we have to remove the normalization applied during local explanation computation to restore the original data. The second metric is Shannon’s entropy shannon1948mathematical , which captures the level of disorder in the heatmap. A lower entropy value suggests more concentrated or structured intensity patterns, whereas a higher entropy value indicates a more uniform distribution of intensity values. The first metric reflects the raw intensity comparison across a series of heatmaps, while the second metric quantifies internal variations within a single heatmap.
We analyzed two XAI methods, SmoothGrad (7(a)) and Grad-CAM++ (7(b)), selecting three mutation configurations guided by the same seed. The first observation is related to gradient attenuation, as the mean intensity of both methods decreases with the confidence, especially those driven by XMutant. Notably, Grad-CAM++’s mean intensity drops significantly, thus it is visualized in a logarithmic scale. Since SmoothGrad reflects gradients across the entire model, the gradient attenuation is less pronounced w.r.t. Grad-CAM++, which instead focuses on local gradients at the last convolutional layer, directly exposing the gradient attenuation in this layer and resulting in a more pronounced decline. Our second observation is related to entropy degradation. The inputs generated by XMutant cause the entropy to increase when the confidence decreases, indicating that the DNN progressively loses its focus. This phenomenon is also more observable with the mutation guided by XMutant equipped with SmoothGrad, since it generates noise-averaged gradients, whereas the entropy of heatmap generated by Grad-CAM++ appears noisy.
5.2 Qualitative Analysis of ADAS
For system-level advanced driving assistance, we considered three configurations of XMutant—Grad-CAM++ (Random), Grad-CAM++ (High)—and further analyzed the driving quality degradation before/after mutation.
Additionally, to allow a fine-grained comparison across approaches for ADAS, we assess the driving quality degradation (DQD) of the lane-keeping model using two metrics 2021-Jahangirova-ICST : (1) the maximum CTE value (cross-track error, i.e., the lateral position of the vehicle w.r.t. the center road) and (2) the Euclidean norm of the steering angle output in the segment of the road affected by the mutation. We record the driving data on the road segment relevant to the candidate control point and derive the driving quality value by applying the selected norm to the raw driving data. Consequently, we calculate DQD by subtracting the pre-mutation driving quality from the post-mutation driving quality.
For each technique, we considered 6,622 pairs of pre- and post-mutation driving quality quantities, involving the maximum CTE value and the Euclidean distance of steering angle in the mutated road segments.
Figure 8: Distributions of Driving Quality Degradation (DQD) based on CTEs/steering angles.
Figure 8 depicts the distribution of DQD, i.e., the subtraction of driving quality pairs. DeepJanus produces only minor variations, with smaller means and a tighter distribution, with its median closely aligned with symmetry around zero. On the other hand, the mutations led by XMutant result in larger mean and variance values for DQD, as well as more data points distributed in areas with high values, indicating a higher driving quality degradation due to more challenging road topologies. Within each approach, we conduct the above-mentioned statistical test over the driving quality distributions before and after mutation. The differences in the quality of driving before/after mutation were found to be statistically significant according to the non-parametric Mann-Whitney U test Wilcoxon1945 ($p$-value $<$ 0.01), which is not the case for the DeepJanus approach.
6 Discussion
XAI for DL Test Generation. In common practice, XAI’s explanations are used qualitatively by humans to understand how a DNN processes the inputs. However, in our research, we use them quantitatively, under the assumption that they contain relevant information related to the behavior of DNNs samek2017explainable ; abs-2201-00009 . Our experiments showed the effectiveness brought by XAI in guiding semantic-based fuzzing, as focusing on the area of attention led to faster discovery of misbehaviors while retaining their validity, as the mutations are more focused and less disruptive.
Our approach depends on the quality of the local explanations, specifically on their capability to constitute a correct reference of the DL system’s behavior. Additionally, understanding the characteristics of these explanations by different XAI methods is crucial for XMutant to leverage them for effective mutation selection and direction. The results of this study showed the significant benefits of utilizing XAI for semantic-based fuzzing. XAI enables the identification of DNN’s misbehaviors while ensuring that the tests remain valid and relevant to the original input domain, thanks to the semantic representation and focused test generation. Although the generation of local explanation introduces additional computational overhead, our findings suggest that the trade-off is justified. Our effectiveness results, even for complex applications such as advanced driving assistance systems, indicate that the benefits of XAI in failure exposure outweigh the costs of computing heatmaps. Our experiments suggest that incorporating XAI (except for the model-agnostic XAI method LIME) into the testing workflow is beneficial. Additionally, the adopted open-source XAI methods can be further optimized for efficiency by parallelizing multiple backpropagations or DNN inferences. We will evaluate this aspect in our future work.
XMutant’s Configurations. All configurations of XMutant are stable in terms of effectiveness across all iterations but our study shows that the choice of the optimal XAI algorithm is task-specific.
For model-level sentiment analysis, both LIME and Integrated Gradients can generate a significant amount of failures, as their explanations differentiate between the positive and negative impacts of semantic concepts, allowing for applying targeted mutations. However, due to LIME’s high computational cost, Integrated Gradients is the most suitable candidate for XMutant in practice for textual inputs. For model-level digit classification, all XAI algorithms are effective, with SmoothGrad (Clustering Attractor) showing the best results. We believe this is associated with its more direct method of reflecting gradient information within the input space, which significantly influences the DNN’s outcome during model-level testing. For system-level automated driving assistance, our findings indicate that Integrated Gradients did not provide benefits compared to the baseline. Conversely, CAM-based algorithms proved highly effective and efficient if a low testing budget was available. However, in system-level testing, it is challenging to identify a reasonable direction for mutation, due to cumulative uncertainty and flakiness of the system 10.1007/s10664-023-10433-5 . This results in targeted mutations (High versus Random) being less effective and offering limited guidance since the system may exhibit different behaviors across consecutive test runs.
XAI-guided Tests preserve validity and original distribution semantics. Thanks to the usage of semantic-based input representation (i.e., a model), XMutant generates failure-inducing inputs while retaining the essential properties of validity and label accuracy. As a result, the generated inputs show a more natural distribution in the embedding space while exhibiting functional novelty, serving as an effective mean to evaluate the model’s generalization ability. In summary, our initial exploration into the usage of XAI for enhancing semantic-based test generation has proved very promising. By carefully selecting and applying the most appropriate XAI methods, it is possible to significantly improve the efficiency and effectiveness of the testing process of complex DL systems.
7 Related Work
With the increasing application of DL to safety-critical domains such as autonomous during, XAI algorithms represent one of the default options to debug the predictions and failures of a DL system. In this section, we focus on the main related propositions.
7.1 Test Generation for Deep Learning Systems
The three main families of DL test generation are semantic-based input representation, raw input manipulation, and latent space manipulation.
Semantic Input Manipulation (SIM) techniques leverage a semantic representation of the input domain (e.g., a model) to generate test inputs, similar to conventional model-driven engineering practices that uphold compliance with domain-specific constraints Abdessalem-ASE18-1 ; Abdessalem-ASE18-2 ; Abdessalem-ICSE18 ; Gambi:2019:ATS:3293882.3330566 ; 2020-Riccio-FSE ; isa . The manipulation occurs on the model, which is subsequently reconverted to the original format larman1998applying . SIM techniques operate within a restricted input space, specifically the control parameters of the model representation. These techniques enhance the realism of the produced outputs by implementing appropriate model constraints.
Several search-based SIM approaches have been applied to DL-based image classifiers. DeepHyperion zohdinasab2021deephyperion uses the MAP-Elites Illumination Search algorithm DBLP:journals/corr/MouretC15 to explore the feature space of the input domain and identify misbehavior-inducing features. DeepMetis 2021-Riccio-ASE a SIM approach that generates inputs that behave correctly on original DL models and misbehave on mutants obtained through injection of realistic faults 2020-Humbatova-ICSE , which can be useful to enhance the mutation killing ability of a test set. DeepJanus 2020-Riccio-FSE is the SIM approach most related to this work since it performs model-based testing of DL systems. Therefore, we performed an explicit empirical comparison with the DeepJanus approach in this work.
To the best of our knowledge, no XAI information is used in existing semantic-based tools to guide test generation. As such, they can be used in conjunction with XMutant to improve their effectiveness.
Raw input manipulation (RIM) techniques involve modifying an image’s original pixel space to create a new input by perturbing the pixel values. RIM techniques aim to produce minimal, often imperceptible changes to original to trigger misbehaviors in the DL system. These methods target different aspects of testing, such as data augmentation or adversarial attacks, which are not directly aligned with our goal. Our method is a functional test generator, differing from adversarial testing in both goals and techniques. Functional testing creates new, valid, in-distribution inputs to evaluate a DNN’s generalization. In contrast, adversarial testing adds minor perturbations to original inputs to test robustness. However, for completeness, we describe the main propositions next.
DeepXplore pei2017deepxplore employs various techniques, including occlusion, light manipulation, and blackout to cause misbehaviors. These perturbations are intended to improve neuron coverage within the DL system. DLFuzz guo2018dlfuzz introduces noise to the seed image to increase the likelihood of system misbehavior. DLFuzz generates adversarial inputs for DL systems without relying on cross-referencing other similar DL systems or manual labeling. DeepTest deeptest alters the images using synthetic affine transformation from the computer vision domain, such as blurring and brightness adjustments, to create simulated rain/fog effects.
Differently, our technique targets functional testing, specifically testing for generalization of DL systems. We achieve this by using XAI guidance in the pixel space, and link it to the model input to generate inputs beyond the original datasets, while remaining within the same distribution. In Section 4, we have showed that our technique generates more natural samples that adhere to the original data’s manifold, whereas these RIM methods, despite their minimal pixel perturbations, artificially inflate failure rates via distribution shifts.
Latent space manipulation (LIM) techniques generate new inputs by learning and reconstructing the underlying distribution of the input data. Sinvad kang2020sinvad constructs the input space using Variational Autoencoders (VAE) kingma2013auto ; doersch2016tutorial and navigates the latent space by adding a random value sampled from a normal distribution to a single element of the latent vector. Sinvad aims to explore the latent space by maximizing either the probability of misbehaviors, estimated from the softmax layer output, or by surprise coverage kim2019guiding . The Feature Perturbations technique dunn2021exposing ; DBLP:journals/corr/abs-2001-11055 involves injecting perturbations into the output of the generative model’s first layers, which represent high-level features of images. These perturbations can affect various characteristics of the image, such as shape, location, texture, or color. DeepRoad deeproad generates driving images using Generative Adversarial Networks (GANs) goodfellow2014generative ; goodfellow2020generative for image-to-image translation. CIT4DNN dola2024cit4dnn combines VAE and combinatorial testing cit605761 . This allows the systematic exploration and generation of diverse and infrequent input datasets. CIT4DNN partitions latent spaces to create test sets that contain a wide range of feature combinations and rare occurrences. A recently proposed technique, Instance Space Analysis, aims to pinpoint the critical features of test scenarios that impact the detection of unsafe behavior isa .
These aforementioned LIM techniques require large-scale training data to capture complete feature distributions, which inherently limits their applicability in resource-constrained or open-world settings. For non-stationary data streams (e.g., evolving NLP token distributions or evolving self-driving scenarios), LIM requires frequent retraining to accommodate changes in the distribution, which incurs excessive computational costs. Indeed, existing testing propositions are limited to DL systems that take as input individual images isa ; dola2024cit4dnn ; deeproad ; 2025-Maryam-ICST ; 2025-Baresi-ICSE . Differently from these LIM approaches, XMutant leverages the local explanation from existing seeds to guide model-based test generation, by perturbing semantic representations instead of latent vectors from a learned manifold. In our paper, we showed that the semantic-based representations made XMutant applicable across different case studies, both at the model and at the system level.
7.2 XAI for DL Testing
Zohdinasab et al. 10304866 compare three state-of-the-art techniques for explanation of DL failures. They show that local and global XAI techniques provide dissimilar explanations for the same inputs and further research is needed to produce better explanations. VisualBackProp 05418 was created to visualize which group of pixels of the input image contributes more to the predictions of a convolutional neural network. Kim and Canny 10631 explore the use of heatmaps for explaining the CNN behavior in an ADS. Xu et al. 09405 investigated the use of XAI techniques to detect action-inducing objects, i.e., objects that have a relevant effect on a driving decision, and jointly predict actions and their respective explanations. ThirdEye 2022-Stocco-ASE focuses on failure prediction of lane-keeping autonomous driving systems during hazardous driving conditions. They turn heatmaps into confidence scores that are used to discriminate safe from unsafe driving behaviors. The intuition is that uncommon heatmaps are associated with unexpected runtime conditions. In this work we use ThirdEye, equipped with the heatmap derivative configuration, to provide such scores during system-level testing.
Fahmy et al. DBLP:journals/corr/abs-2002-00863 apply clustering to heatmaps capturing the relevance of the DNN predictions to automatically support the identification of failure-inducing inputs. Such data is used for the retraining of a gaze detection system that uses DNNs to determine the gaze direction of the driver. The authors present an extension of the previous work arxiv.2204.00480 in which inputs identified by the heatmap-based mechanism are given in input to a search-based test generator.
In contrast, in this work, we use local explanations to support the early detection of corner case inputs of a DL model. In our work we experiment with different XAI algorithms, showing that the choice of the best algorithm is domain and testing level dependent.
8 Conclusions
In this work, we describe and evaluate XMutant, a semantic-based fuzzer for DL systems that generates inputs that focus on the attention of the system under test through mutations that are informed by the local explanation available from XAI algorithms. We evaluated XMutant on both model-level and system-level testing. Our empirical studies show that XMutant is significantly more effective and efficient than the state-of-the-art test generation approaches while preserving a high validity rate of failing test inputs.